# Scalable Message Passing Neural Networks: No Need for Attention in Large Graph Representation Learning

### v1

@@ 4 Theoretical Justification
In this section, we theoretically justify our architectural design. First, we review some theoretical findings from the literature that hint at the importance of residual connections to avoid oversmoothing in message-passing convolutions over graphs; moreover, by the results of [37] they make the loss landscape of the deep learning model much less jagged. Inspired by these results, we further extend the analysis through the lenses of universal approximation.
4.1 Oversmoothing
Oversmoothing is regarded as the tendency of node features to approach the same value after several message-passing transformations. This phenomenon has prompted the adoption of relatively shallow GNNs, as adding many layers has traditionally resulted in node-level features that are too similar to each other and, hence, indistinguishable for downstream learners. This has hindered scalability and led to models with orders of magnitude fewer parameters than, for instance, counterparts in the literature on 2D generative modelling [38] and LLMs [17], which have already adopted architectures with billions of parameters.
In general, standalone message-passing GNNs tend to behave as low-pass filters and may consequently lead to oversmoothing. However, this does not always need to be the case: graph convolution can also magnify high frequencies if it is augmented with residual connections. Following [4], message-passing is characterized as being low- frequency dominant (LFD) or high-frequency dominant (HFD) depending on the asymptotic behavior of the normalized Dirichlet energy of the system as the number of diffusion layers $l\rightarrow\infty$. These lead to oversmoothing and oversharpening, respectively.
Definition 4.1 (Graph Dirichlet Energy of a Message Passing System [39]).
$\mathfrak{E}^{Dir}(\mathbf{X})\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\frac{1}{2}\sum_{i,j\in\mathcal{E}}||(\nabla\mathbf{X})_{ij}||^{2}\,;\,(\nabla\mathbf{X})_{ij}\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\frac{\mathbf{x}_{j}}{\sqrt{\mathrm{deg}(v_{j})}}-\frac{\mathbf{x}_{i}}{\sqrt{\mathrm{deg}(v_{i})}},$ (9)
Definition 4.2 (Graph Frequency Dominance [4]).
We say a message passing GNN is LFD if its normalized Dirichlet energy $\mathfrak{E}^{Dir}(\mathbf{X}^{(l)})/||\mathbf{X}^{(l)}||^{2}\rightarrow\lambda_{1}=0$ for $l\rightarrow\infty$ and HFD if $\mathfrak{E}^{Dir}(\mathbf{X}^{(l)})/||\mathbf{X}^{(l)}||^{2}\rightarrow\lambda_{n}$ for $l\rightarrow\infty$, where $\mathbf{X}^{(l)}$ encapsulates the node level features at a given layer $l$ and $21\leq\lambda_{n}<$ 2 [40] is the largest eigenvalue whose corresponding eigenvector of the normalized graph Laplacian captures fine-grained microscopic behavior of the signal on the graph.
In particular, the above characterization relies on analyzing message-passing from the perspective of gradient flows in the limit of infinitely many layers. Although analyzing the asymptotic behavior of GNNs is interesting from a theoretical perspective, the normalized Dirichlet energy is not directly predictive of model success, since both over-smoothing and over-sharpening will lead to degradation in performance [4]. We provide an alternative perspective through the lens of universal approximation, which does not rely on asymptotic behavior. Specifically, we demonstrate that the universal approximation property of downstream classifiers is compromised when passing input features through graph convolution layers, but it can be restored with a residual connection.
4.2 Universality and Residual Connections
We will use $\mathcal{C}(\mathbb{R}^{N\times D})$ to denote the set of continuous functions from $\mathbb{R}^{N\times D}$ to $\mathbb{R}$, when both are equipped with the (only) norm topologies thereon. We use $\mathcal{MLP}_{N\times D}$ to denote the set of MLPs from $\mathbb{R}^{N\times D}$ to $\mathbb{R}$ with a $\operatorname{SiLU}$ activation function. We consider a class of models to be expressive if it is a universal approximator, in the local uniform sense of [41, 42, 43, 44], meaning that it can approximately implement any continuous function uniformly on compact sets.
@@ 4.2 Universality and Residual Connections
A subset of models $\mathcal{F}\subseteq\mathcal{C}(\mathbb{R}^{N\times D})$ is said to be a universal approximator in $\mathcal{C}(\mathbb{R}^{N\times D})$ if: for each continuous target function $f:\mathbb{R}^{N\times D}\to\mathbb{R}$, each uniform approximation error $\varepsilon>0$ and every non-empty compact set of inputs $K\subset\mathbb{R}^{N\times D}$, there exists some model $\hat{f}\in\mathcal{F}$ such that
@@
We will consider two classes of models, highlighting the key benefits and drawbacks of adding versus omitting residual connections in graph convolution layers (Equation 2). Fix a matrix $\mathbf{W}\in\mathbb{R}^{D\times D}$ and a connected graph $\mathcal{G}$ on $N$ nodes.
No Residual Connection. Consider the class $\mathcal{F}_{\mathcal{G},\mathbf{W}}$ which consists of all maps $\hat{f}:\mathbb{R}^{N\times D}\to\mathbb{R}$
$\hat{f}=\operatorname{MLP}\circ\mathcal{L}_{\mathcal{G},\mathbf{W}}^{\operatorname{conv}}\,;\,\mathcal{L}_{\mathcal{G},\mathbf{W}}^{\operatorname{conv}}(\mathbf{X})\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\tilde{\mathbf{A}}\mathbf{X}\mathbf{W},$ (11)
where $\operatorname{MLP}\in\mathcal{MLP}_{N\times D}$ and $\tilde{\mathbf{A}}$ is the degree-normalized adjacency matrix. The class $\mathcal{F}_{\mathcal{G},\mathbf{W}}$ is the stylized representation of GCN models without a residual connection.
Residual Connection. We benchmark this class against a simplified version of our SMPNN block. Again, fixing $\mathbf{W}$ and a connected graph $\mathcal{G}$ on $N$ nodes. This class, denoted by $\mathcal{F}^{\operatorname{residual}}_{\mathcal{G},\mathbf{W}}$, consists of maps $\hat{f}:\mathbb{R}^{N\times D}\to\mathbb{R}$ with representation
$\hat{f}=\operatorname{MLP}\circ\mathcal{L}_{\mathcal{G},\mathbf{W}}^{\operatorname{conv+r}}\,;\,\hfill\mathcal{L}_{\mathcal{G},\mathbf{W}}^{\operatorname{conv+r}}(\mathbf{X})\stackrel{{\scriptstyle\mbox{\tiny def.}}}{{=}}\tilde{\mathbf{A}}\mathbf{X}\mathbf{W}+\mathbf{X},$ (12)
We emphasize that the MLPs within both sets, $\mathcal{F}_{\mathcal{G},\mathbf{W}}$ and $\mathcal{F}^{\operatorname{residual}}_{\mathcal{G},\mathbf{W}}$, are identical. Hence, there are no inherent advantages or disadvantages between them; the sole discrepancy lies in the presence or absence of a residual connection. The following set of results provides a counter-example to the universality of $\mathcal{F}_{G,\mathbf{W}}$ when $G$ is a complete graph. We shows that this deficit it not an issue when incorporating a residual connection as $\mathcal{F}_{G,\mathbf{W}}^{\text{residual}}$ is universal.
In general, analyzing the expressive power of GNNs from the perspective of universal approximation can be cumbersome due to the variety of vastly different graphs encountered in practice. However, the importance of accompanying graph convolution with a residual connection can already be observed in the case of a complete graph. Our first main finding is a negative result, showing that the universal approximation property of MLPs is lost when passing input features through a graph convolution layer with no residual connection for any learnable weight matrix $\mathbf{W}$ used in equation 11.
@@
Let $N,D$ be positive integers with $N\geq 2$. Let $\mathcal{G}$ be a complete graph on $N$ nodes. For any weight matrix $\mathbf{W}\in\mathbb{R}^{D\times D}$, the class $\mathcal{F}_{\mathcal{G},\mathbf{W}}$ defined in equation 11 is not a universal approximator in $\mathcal{C}(\mathbb{R}^{N\times D})$.
@@
Theorem 4.2 (Universal Approximation via Graph Convolution is Possible with Residual Connections).
Let $N,D$ be positive integers. There exists a weight matrix $\mathbf{W}\in\mathbb{R}^{D\times D}$ such that the class “with skip connection” $\mathcal{F}^{\operatorname{residual}}_{\mathcal{G},\mathbf{W}}$ is a universal approximator in $\mathcal{C}(\mathbb{R}^{N\times D})$.
Interestingly, a more technical version of Theorem 4.2 in the appendix of the paper (Theorem B.6), which shows that the matrix $\mathbf{W}$ allows for universality to be regained through the skip connection is not obscure; rather it can be explicitly constructed by sampling from a random Gaussian matrix. Furthermore such a matrix is nearly always obtained in this manner with probability approaching $100\%$ in high dimensions. All proofs and derivations are provided in Appendix B.
This analysis supports our architectural choices in Section 3.2, which represent an amalgamation of best practices in the LLM literature with the theoretical insights here presented concerning oversmoothing and universality.
@@ 5 Experimental Validation
We experimentally validate our architecture and compare it to recent state-of-the-art baselines. We follow the evaluation protocol used in NodeFormer [10], DIFFormer [11], and SGFormer [12]. Additionally, we perform ablations and further experiments to verify the theoretical analysis presented in Section 4. For details, refer to Appendix C.
Large Scale Graph Datasets (Table 2). We compare the performance of SMPNNs to that of Graph Transformers tailored for large graph transductive learning, as well as other GNN baselines. Note that most Graph Transformer architectures (Table 1) are difficult to scale to these datasets. We observe that SMPNNs consistently outperform SOTA architectures without the need for attention. Augmenting the base SMPNN model with linear attention (Appendix A) leads to improvements in performance of under $1\%$ only, while substantially increasing computational overhead. For instance, in the case of the ogbn-products dataset, using the same hyperparameters, adding linear global attention with a single head increases the total number of model parameters from 834K to 2.4M. The resulting performance gain of only $0.18\%$ requires thus more than twice as many parameters.
Most Graph Transformers in the literature use multiple attention heads (Table 1); we experimented with up to 4 attention heads but found differences in terms of performance not to be statistically significant, which seems to align with the conclusions of [12]. Furthermore, note that as reported in the OGB benchmark paper [2], the Max Strongly Connected Component Ratio (MaxSCC Ratio) for the node-level datasets used in these experiments is 1.00 for all datasets (except for ogbn-products, which is 0.97). This indicates that the entire graph is a single strongly connected component, meaning every node is reachable from every other node. Although this does not guarantee that information bottlenecks will not be present, a high MaxSCC Ratio denotes high inter-connectivity, where information or influence can rapidly spread among a large proportion of the network, which may explain why attention is not that important in such settings. Apart from SMPNN and SMPNN augmented with attention, we also include a Linear Transformer baseline, which corresponds to removing the GCN layer from our architecture and substituting local convolution with global linear attention. This is also clearly worse than SMPNNs, which again verifies that in these datasets, the locality inductive bias is important.
@@
Dataset ogbn-proteins pokec ogbn-arxiv ogbn-products Nodes 132,534 1,632,803 169,343 2,449,029 Edges 39,561,252 30,622,564 1,166,243 61,859,140 MLP 72.04 ± 0.48 60.15 ± 0.03 55.50 ± 0.23 63.46 ± 0.10 GCN [30] 72.51 ± 0.35 62.31 ± 1.13 71.74 ± 0.29 83.90 ± 0.10 SGC [15] 70.31 ± 0.23 52.03 ± 0.84 67.79 ± 0.27 81.21 ± 0.12 GCN-NSampler [30, 45] 73.51 ± 1.31 63.75 ± 0.77 68.50 ± 0.23 83.84 ± 0.42 GAT-NSampler [46, 45] 74.63 ± 1.24 62.32 ± 0.65 67.63 ± 0.23 85.17 ± 0.32 SIGN [16] 71.24 ± 0.46 68.01 ± 0.25 70.28 ± 0.25 80.98 ± 0.31 NodeFormer [10] 77.45 ± 1.15 68.32 ± 0.45 59.90 ± 0.42 87.85 ± 0.24 Difformer [11] 79.49 ± 0.44 69.24 ± 0.76 64.94 ± 0.25 84.00 ± 0.07 SGFormer [12] 79.53 ± 0.38 73.76 ± 0.24 72.63 ± 0.13 89.09 ± 0.10 SMPNN 83.15 ± 0.24 79.76 ± 0.19 73.75 ± 0.24 90.61 ± 0.05 SMPNN + Attention 83.65 ± 0.35 80.09 ± 0.12 74.38 ± 0.16 90.79 ± 0.04 Linear Transformer 60.87 ± 7.23 62.72 ± 0.09 58.02 ± 0.21 79.85 ± 0.08
100M Node Graph Dataset and Scalability Experiments (Table 3). We test the scalability of our pipeline on the ogbn-papers-100M dataset. Our main competitors are SIGN [16] and SGFormer [12], as other Graph Transformers have not been able to scale to this dataset. We outperform SGFormer without requiring attention, again demonstrating the scalability and effectiveness of SMPNNs.
@@
GCN [30] 63.29 ± 0.19
SGC [15] 63.29 ± 0.19
GCN-NSampler [30, 45] 62.04 ± 0.27
GAT-NSampler [46, 45] 63.47 ± 0.39
SIGN [16] 65.11 ± 0.14
SGFormer [12] 66.01 ± 0.37
@@
Additional Image, Text, and Spatio-Temporal Benchmarks (Tables 4 and  5). For completeness, we also test our architecture on image and text classification with low label rates, as well as on spatio-temporal dynamics prediction, following [11]. Our model achieves the best performance in most configurations for the CIFAR, STL, and 20News datasets. Likewise, in Table 5, SMPNNs are also competitive in spatio-temporal prediction. These results demonstrate that our architecture is applicable to a variety of tasks.
@@
Dataset CIFAR STL 20News Labels 100 500 1000 100 500 1000 1000 2000 4000 MLP 65.9 ± 1.3 73.2 ± 0.4 75.4 ± 0.6 66.2 ± 1.4 73.0 ± 0.8 75.0 ± 0.8 54.1 ± 0.9 57.8 ± 0.9 62.4 ± 0.6 ManiReg 67.0 ± 1.9 72.6 ± 1.2 74.3 ± 0.4 66.5 ± 1.9 72.5 ± 0.5 74.2 ± 0.5 56.3 ± 1.2 60.0 ± 0.8 63.6 ± 0.7 GCN-kNN 66.7 ± 1.5 72.9 ± 0.4 74.7 ± 0.5 66.9 ± 0.5 72.1 ± 0.8 73.7 ± 0.4 56.1 ± 0.6 60.6 ± 1.3 64.3 ± 1.0 GAT-kNN 66.0 ± 2.1 72.4 ± 0.5 74.1 ± 0.5 66.5 ± 0.8 72.0 ± 0.8 73.9 ± 0.6 55.2 ± 0.8 59.1 ± 2.2 62.9 ± 0.7 Dense-GAT OOM OOM OOM OOM OOM OOM 54.6 ± 0.2 59.3 ± 1.4 62.4 ± 1.0 GLCN 66.6 ± 1.4 72.8 ± 0.5 74.7 ± 0.3 66.4 ± 0.8 72.4 ± 1.3 74.3 ± 0.7 56.2 ± 0.8 60.2 ± 0.7 64.1 ± 0.8 Difformer-a 69.3 ± 1.4 74.0 ± 0.6 75.9 ± 0.3 66.8 ± 1.1 72.9 ± 0.7 75.3 ± 0.6 57.9 ± 0.7 61.3 ± 1.0 64.8 ± 1.0 Difformer-s 69.1 ± 1.1 74.8 ± 0.5 76.6 ± 0.3 67.8 ± 1.1 73.7 ± 0.6 76.4 ± 0.5 57.7 ± 0.3 61.2 ± 0.6 65.9 ± 0.8 SMPNN 68.6 ± 1.8 76.2 ± 0.5 78.0 ± 0.3 67.9 ± 0.9 73.9 ± 0.7 76.7 ± 0.5 58.9 ± 0.8 62.7 ± 0.6 65.6 ± 0.6
@@
Ablation on SMPNN Architecture Components. In Table 6, we conduct an ablation study. First, we ablate the standard SMPNN architecture by removing residual connections after graph convolution layers. In line with the theory presented in Section 4, we observe that this indeed has a significant impact on model performance. Next, we test the effect of fixing the learnable scaling connections ($\alpha_{1}^{(l)}=\alpha_{2}^{(l)}=1,\forall l$) while retaining the residuals. This decreases performance for ogbn-proteins and ogbn-arxiv, whereas it slightly increases performance for pokec and ogbn-products. We decide to keep the scalings since they make the model more expressive in general. Additionally, we test removing the pointwise feedforward transformation, which leads to a slight drop in performance for all datasets and seems particularly relevant for ogbn-proteins. This experiment suggests that the most critical part of the architecture is the message-passing component, as expected. Lastly, for completeness, we also test removing the LayerNorm before the GCN, which generally leads to a drop in performance. However, in the case of ogbn-arxiv, this configuration obtains a test accuracy of $74.46\%$, which outperforms all other models and baselines. In conclusion, there is no free lunch; depending on the dataset, slight modifications may lead to improved performance and, interestingly, have even more of an effect than augmenting SMPNNs with linear attention. Nevertheless, we find that, in general, the standard SMPNN block is robust and leads to SOTA results across several datasets.
@@
Table 6: Results (mean test set accuracy ± standard deviation over 5 runs) for ablation studies on OGBN large graph datasets.

### v2

@@ 4 Theoretical Justification
In this section, we theoretically justify our architectural design. We review some theoretical findings from the literature that hint at the importance of residual connections to avoid oversmoothing in message-passing convolutions over graphs in Appendix C. Inspired by these results, we further extend the analysis through the lens of universal approximation, which does not rely on asymptotic behavior, unlike other works such as Giovanni et al. (2023). For the proofs, see Appendix D.
We consider a class of models to be expressive if it is a universal approximator, in the local uniform sense of Hornik et al. (1989); Kidger and Lyons (2020); Duan et al. (2023); Yarotsky (2024), meaning that it can approximately implement any continuous function uniformly on compact sets. We will use $\mathcal{C}(\mathbb{R}^{N\times D})$ to denote the set of continuous functions from $\mathbb{R}^{N\times D}$ to $\mathbb{R}$, when both are equipped with the (only) norm topologies thereon.
@@
A subset of models $\mathcal{F}\subseteq\mathcal{C}(\mathbb{R}^{N\times D})$ is said to be a universal approximator in $\mathcal{C}(\mathbb{R}^{N\times D})$ if: for each continuous target function $f:\mathbb{R}^{N\times D}\to\mathbb{R}$, each uniform approximation error $\varepsilon>0$, and every non-empty compact set of inputs $K\subset\mathbb{R}^{N\times D}$, there exists some model $\hat{f}\in\mathcal{F}$ such that
@@
In what follows, for any activation function $\zeta:\mathbb{R}\to\mathbb{R}$ we use $\mathcal{MLP}_{N\times D}^{\zeta}=\mathcal{MLP}_{N\times D}$ to denote the class of MLPs with activation function $\zeta$, mapping $\mathbb{R}^{N\times D}$ to $\mathbb{R}$. We consider the following regularity condition on $\zeta$.
Assumption 4.2 (Regular Activation).
The activation function $\zeta:\mathbb{R}\to\mathbb{R}$ is continuous and either: (i) Pinkus (1999): $\zeta$ is non-polynomial, (ii) Kidger and Lyons (2020): $\zeta$ is non-affine and there exists some $t\in\mathbb{R}$ at which $\zeta$ is continuously differentiable and $\zeta^{\prime}(t)\neq 0$.
We will consider two classes of models, highlighting the key benefits and drawbacks of adding versus omitting residual connections in graph convolution layers.
Definition 4.3 (No Residual Connection Model Class).
Consider the class $\mathcal{F}_{\mathcal{G},\mathbf{W}}$ which consists of all maps $\hat{f}:\mathbb{R}^{N\times D}\to\mathbb{R}$
$\hat{f}=\operatorname{MLP}\circ\mathcal{L}_{\mathcal{G},\mathbf{W}}^{\operatorname{conv}}\,;\,\mathcal{L}_{\mathcal{G},\mathbf{W}}^{\operatorname{conv}}(\mathbf{X})\stackrel{{\scriptstyle\mbox{def.}}}{{=}}\tilde{\mathbf{A}}\mathbf{X}\mathbf{W},$ (6)
where $\mathbf{W}\in\mathbb{R}^{D\times D}$ is a matrix, $\operatorname{MLP}\in\mathcal{MLP}_{N\times D}$ and $\tilde{\mathbf{A}}$ is the degree-normalized adjacency matrix for the input graph $\mathcal{G}$. The class $\mathcal{F}_{\mathcal{G},\mathbf{W}}$ is a stylized model class consisting of a graph convolution layer followed by an MLP, without a residual connection.
We benchmark this class against a simplified version of our SMPNN block.
Definition 4.4 (Residual Connection Model Class).
Again, fix $\mathbf{W}\in\mathbb{R}^{D\times D}$ and a graph $\mathcal{G}$ on $N$ nodes. This class, denoted by $\mathcal{F}^{\operatorname{residual}}_{\mathcal{G},\mathbf{W}}$, consists of maps $\hat{f}:\mathbb{R}^{N\times D}\to\mathbb{R}$ with representation
$\hat{f}=\operatorname{MLP}\circ\mathcal{L}_{\mathcal{G},\mathbf{W}}^{\operatorname{conv+r}}\,;\,\hfill\mathcal{L}_{\mathcal{G},\mathbf{W}}^{\operatorname{conv+r}}(\mathbf{X})\stackrel{{\scriptstyle\mbox{def.}}}{{=}}\tilde{\mathbf{A}}\mathbf{X}\mathbf{W}+\mathbf{X}.$ (7)
We emphasize that the MLPs within both sets, $\mathcal{F}_{\mathcal{G},\mathbf{W}}$ and $\mathcal{F}^{\operatorname{residual}}_{\mathcal{G},\mathbf{W}}$, are identical. Hence, there are no inherent advantages or disadvantages between them; the sole discrepancy lies in the presence or absence of a residual connection.
The following set of results provides a counterexample to the universality of $\mathcal{F}_{\mathcal{G},\mathbf{W}}$ when $\mathcal{G}$ is a complete graph with self-loops (worst-case analysis) and further extends it to general graphs (like those found in practice). We show that this deficit is not an issue when incorporating a residual connection as $\mathcal{F}_{\mathcal{G},\mathbf{W}}^{\text{residual}}$ is universal.
@@
Let $N,D$ be positive integers with $N\geq 2$. Let $\mathcal{G}$ be a complete graph (with self-loops) on $N$ nodes. For any weight matrix $\mathbf{W}\in\mathbb{R}^{D\times D}$, the class $\mathcal{F}_{\mathcal{G},\mathbf{W}}$ defined in equation 6 is not a universal approximator in $\mathcal{C}(\mathbb{R}^{N\times D})$.
@@
Lemma 4.6 (Injectivity criterion on the self-looped complete graph with residuals).
Let $N,D\geq 1$, let $\tilde{\mathbf{A}}=J_{N}/N$, and define $\mathcal{L}_{\mathcal{G},\mathbf{W}}^{\operatorname{conv+r}}(\mathbf{X}):=\tilde{\mathbf{A}}\mathbf{X}\mathbf{W}+\mathbf{X}.$ Then $\mathcal{L}_{\mathcal{G},\mathbf{W}}^{\operatorname{conv+r}}$ is injective if and only if $I_{D}+\mathbf{W}$ is invertible, equivalently $1\notin\sigma(\mathbf{W}).$
Theorem 4.7 (Universality with residual connections).
Let $\zeta$ satisfy Assumption 4.2. Let $N,D\geq 1$, let $\tilde{\mathbf{A}}=J_{N}/N$, and let $\mathbf{W}\in\mathbb{R}^{D\times D}$. If $1\notin\sigma(\mathbf{W})$, then the class $\mathcal{F}^{\operatorname{residual}}_{\mathcal{G},\mathbf{W}}$ is dense in $\mathcal{C}(\mathbb{R}^{N\times D})$ for the topology of uniform convergence on compact sets. In particular, if $\mathbf{W}$ has an absolutely continuous distribution on $\mathbb{R}^{D\times D}$ (for example, i.i.d. Gaussian entries), then this holds with probability one.
The residual connections restore universality of the model class, while the pathological case where injectivity fails occurs with probability zero and is therefore (extremely) unlikely to arise in practice at initialization. Finally, we extend our results to general graphs like those found in experiments.
Corollary 4.8 (Sufficient condition for universality on general graphs).
Let $\mathcal{G}$ be any graph with normalized adjacency matrix $\tilde{\mathbf{A}}\in\mathbb{R}^{N\times N}$, and let $\mathbf{W}\in\mathbb{R}^{D\times D}$. If the spectral norm
$\|\tilde{\mathbf{A}}\|_{2}\,\|\mathbf{W}\|_{2}<1,$
then $\mathcal{L}^{\operatorname{conv+r}}_{\mathcal{G},\mathbf{W}}$ is injective, and therefore the model class $\mathcal{F}^{\operatorname{residual}}_{\mathcal{G},\mathbf{W}}$ is dense in $\mathcal{C}(\mathbb{R}^{N\times D})$.
Thus, in the presence of residual connections, the expressivity of the architecture is linked to the scale of the weights $\mathbf{W}$ relative to the graph connectivity operator $\tilde{\mathbf{A}}$. This could inspire initialization techniques to guarantee that at the onset of training, the model resides in a “universal” regime (which depends on the input graph) where node-level information is preserved rather than collapsed.
@@ 5 Experimental Validation
We experimentally validate our architecture and compare it to recent state-of-the-art baselines. We follow the evaluation protocol used in NodeFormer (Wu et al., 2022), DIFFormer (Wu et al., 2023a), and SGFormer (Wu et al., 2023b). Additionally, we perform ablations and further experiments to verify the theoretical analysis presented in Section 4. For details, refer to Appendix E.
Large Scale Graph Datasets (Table 2). We compare the performance of SMPNNs to that of Graph Transformers tailored for large graph transductive learning, as well as other GNN baselines. Note that most Graph Transformer architectures (Table 1) are difficult to scale to these datasets. We observe that SMPNNs consistently outperform SOTA architectures without the need for attention. Augmenting the base SMPNN model with linear attention (Appendix B) leads to improvements in performance of under $1\%$ only, while substantially increasing computational overhead. For instance, in the case of the ogbn-products dataset, using the same hyperparameters, adding linear global attention with a single head increases the total number of model parameters from 834K to 2.4M. The resulting performance gain of only $0.18\%$ thus requires more than twice as many parameters.
@@
Dataset ogbn-proteins pokec ogbn-arxiv ogbn-products Nodes 132,534 1,632,803 169,343 2,449,029 Edges 39,561,252 30,622,564 1,166,243 61,859,140 MLP 72.04 ± 0.48 60.15 ± 0.03 55.50 ± 0.23 63.46 ± 0.10 GCN (Kipf and Welling, 2017) 72.51 ± 0.35 62.31 ± 1.13 71.74 ± 0.29 83.90 ± 0.10 SGC (Wu et al., 2019) 70.31 ± 0.23 52.03 ± 0.84 67.79 ± 0.27 81.21 ± 0.12 GCN-NSampler Kipf and Welling (2017); Zeng et al. (2020) 73.51 ± 1.31 63.75 ± 0.77 68.50 ± 0.23 83.84 ± 0.42 GAT-NSampler Veličković et al. (2018); Zeng et al. (2020) 74.63 ± 1.24 62.32 ± 0.65 67.63 ± 0.23 85.17 ± 0.32 SIGN (Frasca et al., 2020) 71.24 ± 0.46 68.01 ± 0.25 70.28 ± 0.25 80.98 ± 0.31 NodeFormer (Wu et al., 2022) 77.45 ± 1.15 68.32 ± 0.45 59.90 ± 0.42 87.85 ± 0.24 DIFFormer (Wu et al., 2023a) 79.49 ± 0.44 69.24 ± 0.76 64.94 ± 0.25 84.00 ± 0.07 SGFormer (Wu et al., 2023b) 79.53 ± 0.38 73.76 ± 0.24 72.63 ± 0.13 89.09 ± 0.10 SMPNN 83.15 ± 0.24 79.76 ± 0.19 73.75 ± 0.24 90.61 ± 0.05 SMPNN + Attention 83.65 ± 0.35 80.09 ± 0.12 74.38 ± 0.16 90.79 ± 0.04 Linear Transformer 60.87 ± 7.23 62.72 ± 0.09 58.02 ± 0.21 79.85 ± 0.08
Most Graph Transformers in the literature use multiple attention heads (Table 1); we experimented with up to 4 attention heads but found differences in terms of performance not to be statistically significant, which seems to align with the conclusions of (Wu et al., 2023b). Furthermore, note that as reported in the OGB benchmark paper (Hu et al., 2020), the Max Strongly Connected Component Ratio (MaxSCC Ratio) for the node-level datasets used in these experiments is 1.00 for all datasets (except for ogbn-products, which is 0.97). This indicates that the entire graph is a single strongly connected component, meaning every node is reachable from every other node. Although this does not guarantee that information bottlenecks will not be present, a high MaxSCC Ratio denotes high inter-connectivity, where information or influence can rapidly spread among a large proportion of the network, which may explain why attention is not that important in such settings. Apart from SMPNN and SMPNN augmented with attention, we also include a Linear Transformer baseline, which corresponds to removing the GCN layer from our architecture and substituting local convolution with global linear attention. This also performs worse than SMPNNs, which suggests that in these datasets, the locality inductive bias is important.
100M Node Graph Dataset and Scalability Experiments (Table 3). We test the scalability of our pipeline on the ogbn-papers-100M dataset. Our main competitors are SIGN (Frasca et al., 2020) and SGFormer (Wu et al., 2023b), as other Graph Transformers have not been able to scale to this dataset. We outperform SGFormer without requiring attention, again demonstrating scalability.
@@
GCN (Kipf and Welling, 2017) 63.29 ± 0.19
SGC (Wu et al., 2019) 63.29 ± 0.19
GCN-NSampler Kipf and Welling (2017); Zeng et al. (2020) 62.04 ± 0.27
GAT-NSampler Veličković et al. (2018); Zeng et al. (2020) 63.47 ± 0.39
SIGN (Frasca et al., 2020) 65.11 ± 0.14
SGFormer (Wu et al., 2023b) 66.01 ± 0.37
@@
Additional Image, Text, and Spatio-Temporal Benchmarks (Tables 4 and  5). For completeness, we also test our architecture on image and text classification with low label rates, as well as on spatio-temporal dynamics prediction, following (Wu et al., 2023a). Our model achieves the best performance in most configurations for the CIFAR, STL, and 20News datasets. Likewise, in Table 5, SMPNNs are also competitive in spatio-temporal prediction. These results demonstrate that our architecture is applicable to a variety of tasks.
@@
Dataset CIFAR STL 20News Labels 100 500 1000 100 500 1000 1000 2000 4000 MLP 65.9 ± 1.3 73.2 ± 0.4 75.4 ± 0.6 66.2 ± 1.4 73.0 ± 0.8 75.0 ± 0.8 54.1 ± 0.9 57.8 ± 0.9 62.4 ± 0.6 ManiReg 67.0 ± 1.9 72.6 ± 1.2 74.3 ± 0.4 66.5 ± 1.9 72.5 ± 0.5 74.2 ± 0.5 56.3 ± 1.2 60.0 ± 0.8 63.6 ± 0.7 GCN-kNN 66.7 ± 1.5 72.9 ± 0.4 74.7 ± 0.5 66.9 ± 0.5 72.1 ± 0.8 73.7 ± 0.4 56.1 ± 0.6 60.6 ± 1.3 64.3 ± 1.0 GAT-kNN 66.0 ± 2.1 72.4 ± 0.5 74.1 ± 0.5 66.5 ± 0.8 72.0 ± 0.8 73.9 ± 0.6 55.2 ± 0.8 59.1 ± 2.2 62.9 ± 0.7 DenseGAT OOM OOM OOM OOM OOM OOM 54.6 ± 0.2 59.3 ± 1.4 62.4 ± 1.0 GLCN 66.6 ± 1.4 72.8 ± 0.5 74.7 ± 0.3 66.4 ± 0.8 72.4 ± 1.3 74.3 ± 0.7 56.2 ± 0.8 60.2 ± 0.7 64.1 ± 0.8 DIFFormer-a 69.3 ± 1.4 74.0 ± 0.6 75.9 ± 0.3 66.8 ± 1.1 72.9 ± 0.7 75.3 ± 0.6 57.9 ± 0.7 61.3 ± 1.0 64.8 ± 1.0 DIFFormer-s 69.1 ± 1.1 74.8 ± 0.5 76.6 ± 0.3 67.8 ± 1.1 73.7 ± 0.6 76.4 ± 0.5 57.7 ± 0.3 61.2 ± 0.6 65.9 ± 0.8 SMPNN 68.6 ± 1.8 76.2 ± 0.5 78.0 ± 0.3 67.9 ± 0.9 73.9 ± 0.7 76.7 ± 0.5 58.9 ± 0.8 62.7 ± 0.6 65.6 ± 0.6
@@
Ablation on SMPNN Architecture Components. In Table 6, we conduct an ablation study. First, we ablate the standard SMPNN architecture by removing residual connections after graph convolution layers. In line with the theory presented in Section 4, we observe that this indeed has a significant impact on model performance. Next, we test the effect of fixing the learnable scaling coefficients ($\alpha_{1}^{(l)}=\alpha_{2}^{(l)}=1,\forall l$) while retaining the residuals. This decreases performance for ogbn-proteins and ogbn-arxiv, whereas it slightly increases performance for pokec and ogbn-products. We decide to keep the scalings since they make the model more expressive in general. Additionally, we test removing the pointwise feedforward transformation, which leads to a slight drop in performance for all datasets and seems particularly relevant for ogbn-proteins. This experiment suggests that the most critical part of the architecture is the message-passing component, as expected. Lastly, for completeness, we also test removing the LayerNorm before the GCN, which generally leads to a drop in performance. However, in the case of ogbn-arxiv, this configuration obtains a test accuracy of $74.46\%$, which outperforms all other models and baselines. In conclusion, there is no free lunch; depending on the dataset, slight modifications may lead to improved performance and, interestingly, have even more of an effect than augmenting SMPNNs with linear attention. Nevertheless, we find that, in general, the standard SMPNN block is robust and leads to SOTA results across several datasets.
@@
Table 6: Results (mean test set accuracy ± standard deviation over 5 runs) for ablation studies on OGBN and pokec large graph datasets.